# Interpreting Graph Inference with Skyline Explanations

### v2

@@ I Introduction
[…]
@@
$\mathsf{fac}$ factual ${\mathcal{M}}(v,G)$ = ${\mathcal{M}}(v,G_{\zeta})?$$\mathsf{true}$,$\mathsf{false}$ a Boolean function  [6, 10, 11, 8]
$\mathsf{cfac}$ counterfactual ${\mathcal{M}}(v,G)\neq{\mathcal{M}}(v,G\setminus G_{\zeta})?$$\mathsf{true}$,$\mathsf{false}$ a Boolean function  [7, 11, 8]
${\mathsf{fdl}}^{+}$ fidelity+ $Pr(M(v,G))-Pr(M(v,G\setminus G_{\zeta}))$ $[-1,1]$ the larger, the better  [14, 12]
${\mathsf{fdl}}^{-}$ fidelity- $Pr(M(v,G))-Pr(M(v,G_{\zeta}))$ $[-1,1]$ the smaller, the better  [14, 10]
@@
$\mathsf{shapley}$ Shapley value $\phi(G_{\zeta})=\sum_{S\subseteq P\setminus\{G_{\zeta}\}}\frac{|S|!(|P|-|S|-1)!}{|P|!}m(S,G_{\zeta})$ $[-1,1]$ total contribution of nodes in $G_{\zeta}$  [12]
@@
Contribution. We summarize our contribution as follows.
[…]
(2) We introduce efficient algorithms to evaluate $\mathsf{SXQs}$. (i) Our first algorithm adopts an “onion-peeling” strategy to iteratively reduce edges at each hop of targeted nodes, and validates a bounded number of generated explanations in a discretized coordination system, to incrementally improve the explainability of the subgraphs. We show that this process ensures a $(1+\epsilon)$ approximation of the optimal solution. (ii) We also present an algorithm to diversify the answer set for $\mathsf{SXQs}$, to provide a comprehensive solution in terms of node embedding and structural differences. (iii) Additionally, we propose parallel algorithms that scale to large graphs by clustering workloads to optimize edge information sharing and load balancing.
[…]
@@
Explanation of GNNs. Several $\mathsf{GNN}$ explanation approaches have been studied. (1) Learning-based methods aim to learn substructures of underlying graphs that contribute to the output of a $\mathsf{GNN}$. GNNExplainer [6] identifies subgraphs with node features that maximize the influence on the prediction by learning continuous soft masks for both adjacency matrix and feature matrix. CF-GNNExplainer [7] learns the counterfactual subgraphs, which lead to significant changes of the output if removed from the graphs. PGExplainer [10] parameterizes the learning process of mask matrix using a multi-layer perceptron. GraphMask [32] learns to mask the edges through each layer of $\mathsf{GNNs}$ that leads to the most sensitive changes to their output. (2) Learning-based $\mathsf{GNN}$ explainers require prior knowledge of model parameters and incur learning overhead for large graphs. Post-hoc approaches perform post-processing to directly compute subgraphs that optimize a fixed criteria. For example,  [12] utilizes Monte-Carlo tree search to compute subgraphs that optimize a game-theory-inspired Shapley value.  [33] follows a similar approach, yet optimizes HN values, a topology-aware variant of Shapley value. These methods are constrained to pre-defined criterion, and cannot easily adapt to customizable, multiple explanability criteria.
@@ II Graphs and GNN Explanation
$\vspace{-1ex}\mathbf{H}^{(l+1)}=\sigma(\widetilde{\mathbf{A}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$ (1)
@@
[…]
@@
$G=(V,E)$ a graph $G$ with node set $V$ and edge set $E$
X, $\widetilde{\mathbf{A}}$ X: feature matrix, $\widetilde{\mathbf{A}}$: normalized adjacency matrix
@@
$H^{(l)}(v,G)$; $H(v,G)$ embedding of node $v$ in $G$ at layer $l$ $(l\in[1,L])$
${\mathcal{M}}(G)$, ${\mathcal{M}}(v,G))$ task-specific output of ${\mathcal{M}}$ over $G$ and over $v\in V$
$v_{t}$; $V_{T}$ a (test) node; a set of (test) nodes
$s$; $G_{s}$; $G_{\zeta}$ a state $s$; a candidate $G_{s}$; an explanatory subgraph
[…]
Several $\mathsf{GNN}$ explainers and the explanability measures they aim to optimize are summarize in Table I. We summarize the main notations in Table II.
@@ III-A Skyline Explanatory Query
We start with a class of explanatory queries. A Skyline explanatory query, denoted as $\mathsf{SXQ}$, has a form ${\mathsf{SXQ}}(G,{\mathcal{M}},v_{t},\Phi)$, where (1) $G$ = $(V,E)$ is an input graph, (2) ${\mathcal{M}}$ is a $\mathsf{GNN}$; (3) $v_{t}\subseteq V_{T}$ is a designated test node of interest, with an output ${\mathcal{M}}(v_{t},G)$ to be explained; and (4) $\Phi$, the measurement space, refers to a set of normalized explanability measures to be maximized, each has a range $(0,1]$. For a measure to be better minimized (e.g., ${\mathsf{conc}}$, ${\mathsf{fdl}}^{-}$ in Table I), one can readily convert it to an inverse counterpart.
To characterize the query semantic, we specify explanatory subgraphs and their dominance relation.
@@
    for any other $G\in\zeta\setminus{\mathcal{G}}_{\zeta}$, and any $G^{\prime}\in{\mathcal{G}}_{\zeta}$, $G\prec G^{\prime}$.
@@
A skyline explanation may contain an excessive number of subgraphs that are too many to inspect. We thus pose a pragmatic cardinality constraint $k$. We say a skyline explanation is a $k$-explanation, if it contains at most $k$ explanatory subgraphs ($k\leq|\zeta|]$). A $k$-skyline query (denoted as ${\mathsf{SXQ}}^{k}$) admits only $k$-explanations as its query answers.
@@ III-C Computational Complexity
Proof sketch:  It suffices to show that ${\mathsf{EVAL}}({\mathsf{SXQ}}^{k})$ for an interpretable space $|\zeta|$ = $f(|G|)$ for some polynomial function $f$, and $|\Phi|=3$. We show this by constructing a polynomial-time reduction from $k$-representative skyline selection problem ($k$-RSP). Given a set $S$ of $d$-dimensional data points, $k$-RSP is to compute a subset $S^{*}$ of $S$, such that (a) $S^{*}$ is a set of skyline points, $|S^{*}|=k$, and (b) $S^{*}$ maximizes the dominance score. $k$-RSP is $\mathsf{NP}$-hard even for $3$-dimensional data points [22].
[…]
@@
A straightforward evaluation of ${\mathsf{SXQ}}^{k}$ may explore the $L$-hop subgraph $G^{L}(\{v_{t}\})$ and initializes the interpretable space $\zeta$ as all connected subgraphs in $G^{L}(\{v_{t}\})$, by invoking subgraph enumeration algorithms [39, 40]. It then enumerates $n$ size-$k$ subsets of $\zeta$ and finds an optimal $k$-explanation. Although this correctly finds optimal explanations, it is not practical for large $G$, as $n$ alone can be already $2^{deg^{L}}$ and the Pareto sets to be inspected can be $\binom{n}{k}$. We thus consider “approximate query processing” scheme for ${\mathsf{SXQ}}^{k}$, and present efficient algorithms that do not require such enumeration.
@@ IV-A Approximating Skyline Explanations
[…]
@@
[…]
@@
Auxiliary structures. $\mathsf{ASX}$-$\mathsf{OP}$ coordinates the interaction of the $\mathsf{Generator}$, $\mathsf{Verifier}$, and $\mathsf{Updater}$ via a state graph (simply denoted as $\zeta^{\prime}$). Each node (a “state”) $s\in\zeta^{\prime}$ records a candidate $G_{s}$ and its local information to be updated and used for evaluating ${\mathsf{SXQ}}^{k}$. There is a directed edge (a “transition”) $t$ = $(s,s^{\prime})$ in $\zeta^{\prime}$ if $G_{s^{\prime}}$ is obtained by applying a graph editing operator (e.g., edge insertion, edge deletion) to $G_{s}$. A path $\rho$ in the state graph $\zeta^{\prime}$ consists of a sequence of transitions that results in a candidate. In addition, each state $s$ is associated with (1) a score ${\sf DS}(G_{s})$, (2) a coordinate $\Phi(s)$, where each entry records an explainability measure $\phi(G_{s})$ ($\phi\in\Phi$); and (3) a variable-length bitvector $B(s)$, where an entry $B(s)[i]$ is $1$ if $G_{i}\preceq_{\epsilon}G_{s}$, and 00 otherwise. The vector $B(s)$ keeps the $\epsilon$-dominance relation between $G_{s}$ and current candidates in $\zeta^{\prime}$. Its ${\sf DS}(G_{s})$ score over $\zeta^{\prime}$, can be readily counted as the number of “1” entries in $B(s)$.
@@
3:set ${\mathcal{G}}_{\epsilon}$:$\emptyset$;
4:identify edges for each hop: ${\mathcal{E}}$:=$\{E_{L},E_{L-1},\ldots,E_{1}\}$;
5:for $l$ = $L$ to $1$ do
6:     initializes state $s_{0}$:=$G^{l}(v_{t})$.
@@
13:               ${\mathcal{G}}_{\epsilon}$:=updateSX $(s,G_{s},\zeta^{\prime},{\mathcal{G}}_{\epsilon})$; $E_{l}$:=$E_{l}\backslash\{e\}$;
@@
8:     if $|{\mathcal{G}}_{\epsilon}|$$<$$k$ then
@@
Update Dominance Relations. Procedure updateSX maintains a set of skyline explanations ${\mathcal{G}}_{s}$ (not shown) in terms of $(1+\epsilon)$-dominance. The set ${\mathcal{G}}_{s}$ is efficiently derived from the bitvectors $B(s)$ of the states in $\zeta^{\prime}$. The latter compactly encodes a lattice structure of $(1+\epsilon)$ dominance as a directed acyclic graph; and ${\mathcal{G}}_{s}$ refers to the states with no “parent” in the lattice, i.e., having an explanatory subgraph that is not $(1+\epsilon)$-dominated by any others so far. Details are in Appendix.
@@
Figure 2: Illustration of Onion Peeling ($L$=$2$, $k$=$2$). $\zeta^{\prime}$ = $\{s_{1},s_{2},s_{3},s_{5}\}$. $(\zeta^{\prime},\epsilon)$-explanation ${\mathcal{G}}_{\epsilon}$=$\{s_{2},s_{5}\}$ with ${\sf DS}$ = $4$.
@@
Proof sketch:  Consider procedure updateSX upon the arrival, at any time, of a new verified candidate $G_{s}$. (1) The above results clearly hold when $|\zeta^{\prime}|\leq k$ or $|{\mathcal{G}}_{\epsilon}|\leq k$, as ${\mathcal{G}}_{\epsilon}$ is the only $(\zeta^{\prime},\epsilon)$-explanation so far. (2) When $|{\mathcal{G}}_{\epsilon}|=k$, we reduce the approximate evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the online MAX $k$-SET COVERAGE problem [43]. The problem maintains a size-$k$ set cover with maximized weights. We show that updateSX adopts a greedy replacement policy by replacing a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a $(1+\frac{1}{k})$ factor improvement for ${\sf DS}$. This ensures a $\frac{1}{4}$-approximation ratio [43]. Following the replacement policy consistently,  updateSX ensures a $\frac{1}{4}$ approximation ratio for $(\zeta^{\prime},\epsilon)$-explanations in $\zeta^{\prime}$. $\Box$
@@
Edge Prioritization. $\mathsf{ASX}$-$\mathsf{OP}$ adopts an edge prioritization heuristic to favor promising candidates with small loss of ${\sf DS}$ and that are more likely to be a skyline explanation ($\mathsf{ASX}$-$\mathsf{OP}$, line 9). It ranks each transaction $t$=$(s,s^{\prime})$ in $\zeta$ based on a loss estimation of ${\sf DS}$ by estimating $\Phi(s^{\prime})$. The cost vector of each transaction is aggregated to an average weight based on each dimension, i.e., $w(t)$ = $\frac{1}{|\Phi|}$ $\sum_{i=1}^{|\Phi|}(\phi_{i}(s)-\hat{\phi_{i}(s^{\prime})})$. The candidates from spawned states with the smallest loss of ${\sf DS}$ are preferred. This helps early convergence of high-quality explanations, and also promotes early detection of non-dominating candidates (see “early pruning”).

### v3

@@ I Introduction
Graph models such as graph neural networks ($\mathsf{GNNs}$) have been trained and routinely queried to perform network analysis in, e.g., biochemistry, social and financial networks [1, 2, 3, 4], among other graph analytical tasks. Given a graph $G$ and a set of test nodes $V_{T}$ in $G$, a $\mathsf{GNN}$ $\mathcal{M}$} can be considered as a function that converts a feature representation of $G$ in the form of $(X,A)$, where $X$ (resp. $A$) refers to the node feature matrix (resp. normalized adjacency matrix) of $G$, to an output representation $Z$ (“output embedding matrix”). The output $Z$ can be post-processed to task-specific output such as labels for classification. An “inference query” invokes the above process to query $\mathcal{M}$}, to perform the inference analysis of $G$.
@@
$\mathsf{fac}$ factual ${\mathcal{M}}(v,G)$ = ${\mathcal{M}}(v,G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [6, 10, 11, 8]
$\mathsf{cfac}$ counterfactual ${\mathcal{M}}(v,G)\neq{\mathcal{M}}(v,G\setminus G_{\zeta})?$ $\mathsf{true}$$\mathsf{false}$ } a Boolean function  [7, 11, 8]
$\mathsf{fdl}$$^{+}$ fidelity+ $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G\setminus G_{\zeta}))$ [1,1][-1,1] the larger, the better  [14, 12]
$\mathsf{fdl}$$^{-}$ fidelity- $Pr({\mathcal{M}}(v,G))-Pr({\mathcal{M}}(v,G_{\zeta}))$ [1,1][-1,1] the smaller, the better  [14, 10]
@@
$\mathsf{shapley}$ Shapley value $\phi(G_{\zeta})=\sum_{S\subseteq P\setminus\{G_{\zeta}\}}\frac{|S|!(|P|-|S|-1)!}{|P|!}m(S,G_{\zeta})$ [1,1][-1,1] total contribution of nodes in $G_{\zeta}$  [12]
@@
Contributions. We summarize our contributions as follows:
(1) Formulation of Skyline Explanations. We introduce a class of skyline explanatory queries ($\mathsf{SXQ}$) that returns a skyline set of explanatory subgraphs optimizing multiple explainability measures. We formalize subgraph dominance and prove the computational hardness of query evaluation problem.
(2) Algorithmic Solutions. We develop efficient algorithms for evaluating $\mathsf{SXQs}$: an onion-peeling method with $(1+\epsilon)$-approximation, a diversification strategy to capture structural and embedding-level variety, and parallel algorithms that scale well to e.g., billion-edge graphs.
(3) Extensive Empirical Evaluation. Experiments on diverse real-world graphs show that our method generates more comprehensive explanations and scales effectively. For example, on $\mathsf{Cora}$(citation network), it improves the Integrated Preference score of state-of-the-art explainers by $2.8\times$, on and achieves an $8.2\times$ speedup over $\mathsf{GExp}$ on the million-edge $\mathsf{OGBN\_arxiv}$ (citation network).
@@
Explanation of GNNs. Several $\mathsf{GNN}$ explanation approaches have been studied. (1) Learning-based methods aim to learn substructures of underlying graphs that contribute to the output of a $\mathsf{GNN}$. GNNExplainer [6] identifies subgraphs with node features that maximize the influence on the prediction by learning continuous soft masks for both adjacency matrix and feature matrix. CF-GNNExplainer [7] learns the counterfactual subgraphs, which lead to significant changes of the output if removed from the graphs. PGExplainer [10] parameterizes the learning process of mask matrix using a multi-layer perceptron. GraphMask [32] learns to mask the edges through each layer of $\mathsf{GNNs}$ that leads to the most sensitive changes to their output. (2) Learning-based $\mathsf{GNN}$ explainers require prior knowledge of model parameters and incur learning overhead for large graphs. Post-hoc approaches perform post-processing to directly compute subgraphs that optimize a fixed criteria. For example,  [12] utilizes Monte-Carlo tree search to compute subgraphs that optimize a game-theory-inspired Shapley value.  [33] follows a similar approach, yet optimizes HN values, a topology-aware variant of Shapley value. SAME [34] extends structure-aware Shapley explanations by using expansion-based Monte Carlo tree search to extract multi-grained connected substructures. These methods are constrained to pre-defined criterion, and cannot easily adapt to customizable, multiple explanability criteria. Beyond post-hoc explainers, GSAT [35] uses stochastic attention within the information bottleneck framework to filter irrelevant components, while PGIB [36] combines prototype learning with the bottleneck to highlight prototypical subgraphs.
@@ II Graphs and GNN Explanation
$\vskip-3.87498pt\mathbf{H}^{(l+1)}=\sigma(\widetilde{\mathbf{A}}\mathbf{H}^{(l)}\mathbf{W}^{(l)})$ (1)
@@
Fixed and Deterministic Inference. We say that a $\mathsf{GNN}$ $\mathcal{M}$} has a fixed inference process if its inference process is specified by fixed model parameters, number of layers, and message passing scheme. It has a deterministic inference process if ${\mathcal{M}}(\cdot)$ generates the same result for the same input. We consider $\mathsf{GNNs}$ with fixed, deterministic inference processes for consistent and robust performance in practice.
@@
GG=$(V,E)$ a graph $G$ with node set $V$ and edge set $E$
@@
${\mathcal{M}}(v,G)$ task-specific output of $\mathcal{M}$} over $v\in V$
$V_{T}$; $v_{t}$; $s$ a set of (test) nodes; $v_{t}\in V_{T}$; a state $s$ (subgraph)
${\mathcal{G}}_{\zeta}$; $G_{\zeta}$; $\zeta$ a set of explanatory subgraphs; $G_{\zeta}\in{\mathcal{G}}_{\zeta}$; interpretable space
$\Phi$; $\phi$ a set of explainability measures; $\phi\in\Phi$
$k$; $\alpha$; $\epsilon$ size of skyline set; dominance score approximation; dominance factor
Node Classification. Node classification is a fundamental task in graph analysis [40]. A $\mathsf{GNN}$-based node classifier learns a $\mathsf{GNN}$ ${\mathcal{M}}:(X,\mathbf{A})\rightarrow\mathbf{Y}$ s.t. ${\mathcal{M}}(v)=y_{v}$ for $v\in V_{Tr}\subseteq V$, where $V_{Tr}$ is the training set of nodes with known (true) labels $Y_{Tr}$. The inference process of a trained $\mathsf{GNN}$ $\mathcal{M}$} assigns the label for a test node $v_{t}\subseteq V_{T}$, which are derived from their computed embeddings.
GNN Explainers and Measures. Given a $\mathsf{GNN}$ $\mathcal{M}$} and an output ${\mathcal{M}}(v,G)$ to be explained, an explanatory subgraph $G_{\zeta}$ is an edge-induced, connected subgraph of $G$ with a non-empty edge set $E_{\zeta}\subseteq E$ that are responsible to clarify the occurrence of ${\mathcal{M}}(v,G)$. We call the set of all explanatory subgraphs as an interpretable space, denoted as $\zeta$.
A $\mathsf{GNN}$ explainer is an algorithm that generates explanatory subgraphs in $\zeta$ for ${\mathcal{M}}(v,G)$. An explainability measure $\phi$ is a function: $G_{\zeta}\rightarrow$$\mathbf{R}$ that associates an explanatory subgraph with an explainability score. Given $G$ and ${\mathcal{M}}(v,G)$ to be explained, existing explainers typically solve a single-objective optimization problem: $G_{\zeta}^{*}$ = $\operatorname*{arg\,max}_{G_{\zeta}\in\zeta}$ $\phi(G_{\zeta})$.
Several $\mathsf{GNN}$ explainers and the explanability measures that they aim to optimize, are summarized in Table I. We summarize the main notations in Table II.
Skyline Query. A skyline query is a multi-criteria database query that retrieves a set of non-dominated records from a dataset, based on a given set of attributes [41, 42]. A record dominates another if it is at least as good in all dimensions and strictly better in at least one. For example, when searching for hotels, a skyline query would return those that are not worse than any other option in both price and distance to the city center, while being strictly better in at least one of them. The result forms the Pareto frontier of the dataset, representing the most interesting or optimal choices without requiring users to specify explicit weighting functions across multiple criteria. Skyline queries are widely used in decision making, recommender systems, and multi-objective optimization tasks.
@@ III-A Skyline Explanatory Query
We start with a class of explanatory queries. A Skyline explanatory query, denoted as $\mathsf{SXQ}$, has a form ${\mathsf{SXQ}}(G,{\mathcal{M}},v_{t},\Phi)$, where (1) $G$ = $(V,E)$ is an input graph, (2) $\mathcal{M}$} is a $\mathsf{GNN}$; (3) $v_{t}\subseteq V_{T}$ is a designated test node of interest, with an output ${\mathcal{M}}(v_{t},G)$ to be explained; and (4) $\Phi$, the measurement space, refers to a set of normalized explainability measures to be maximized, each has a range $(0,1]$. For a measure to be better minimized (e.g., $\mathsf{conc}$}, $\mathsf{fdl}$$^{-}$ in Table I), one can readily convert it to an inverse counterpart. To characterize the query semantic, we specify explanatory subgraphs and their dominance relation.
@@
    for any other $G\in\zeta\setminus{\mathcal{G}}_{\zeta}$, and any $^{\prime}\in{\mathcal{G}}_{\zeta}, GG$$\prec G^{\prime}$.
@@
A skyline explanation may contain an excessive number of subgraphs that are too many to inspect. We thus pose a pragmatic cardinality constraint $k$. We say that a skyline explanation is a $k$-explanation, if it contains at most $k$ explanatory subgraphs ($k\leq|\zeta|]$). A $k$-skyline query (denoted as ${\mathsf{SXQ}}^{k}$) admits only $k$-explanations as its query answers.
@@ III-C Computational Complexity
Proof sketch: It suffices to show that ${\mathsf{EVAL}}({\mathsf{SXQ}}^{k})$ is $\mathsf{NP}$-hard for an interpretable space $|\zeta|$ = $f(|G|)$ for some polynomial function $f$, and $|\Phi|=3$. We show this by constructing a polynomial-time reduction from $k$-representative skyline selection problem ($k$-RSP). Given a set $S$ of $d$-dimensional data points, $k$-RSP is to compute a subset $S^{*}$ of $S$, such that (a) $S^{*}$ is a set of skyline points, $|S^{*}|=k$, and (b) $S^{*}$ maximizes the dominance score. $k$-RSP is $\mathsf{NP}$-hard even for $3$-dimensional data points [22].
Given an instance of $k$-RSP with a set $S$ of $n$ 33-dimensional data points, the reduction sets a graph $G$ as a two-level tree with a single root $v_{t}$, and $n$ distinct edges $(v_{i},v_{t})$, one for each data point $s_{i}\in S$. All the nodes in $G$ are assigned the same feature vector and the same label $l$. Duplicate graph $G$ as a training graph $G_{T}$ and train a vanilla $\mathsf{GCN}$ $\mathcal{M}$} with LL=22, with an output layer a trivial function that sets an output ${\mathcal{M}}(v_{t},G)$ = $l$. As such, any sub-tree rooted at $v_{t}$ in $G$ is a factual explanatory subgraph $G_{\epsilon}$ for the output ${\mathcal{M}}(v_{t},G)$ = $l$. For $\Phi$, we fix one of the measures $\phi_{1}\in\Phi$ to be “conciseness”, hence enforcing the interpretable space to contain only singleton edges, i.e., any explanatory subgraph $G_{\epsilon}^{\prime}$ with more than one edges is strictly dominated by a single edge counterpart.
@@
A straightforward evaluation of ${\mathsf{SXQ}}^{k}$ may explore the $L$-hop subgraph $G^{L}(\{v_{t}\})$ and initializes the interpretable space $\zeta$ as all connected subgraphs in $G^{L}(\{v_{t}\})$, by invoking subgraph enumeration algorithms [44, 45]. It then enumerates $n$ size-$k$ subsets of $\zeta$ and finds an optimal $k$-explanation. Although this correctly finds optimal explanations, it is not practical for large $G$, as $n$ alone can be $2^{deg^{L}}$ and the Pareto sets inspection can take $O(\binom{n}{k})$ time. We thus consider “approximate query processing” scheme for ${\mathsf{SXQ}}^{k}$, and present efficient algorithms that do not require such enumeration.
@@ IV-A Approximating Skyline Explanations
Here (1) $r_{\Phi}$ = $\max\frac{\phi_{u}}{\phi_{l}}$, for each measure $\phi\in\Phi$ with a range $\phi_{l},\phi_{u}]$; (2) $G^{L}(v_{t})$ refers to the $L$-hop neighbor subgraph of node $v_{t}$, and (3) $L$ is the number of layers of the $\mathsf{GNN}$ $\mathcal{M}$}. Note that in practice, $|\Phi|$, $L$, and $\epsilon\in[0,1]$ are small constants, and $r_{\Phi}$ is often small.
@@
Algorithm. Our first algorithm, denoted as $\mathsf{ASX}$-$\mathsf{OP}$ (illustrated as Algorithm 1), takes advantage of a data locality property: For a $\mathsf{GNN}$ $\mathcal{M}$} with $L$ layers and any node $v_{t}$ in $G$, its inference only involves the nodes up to $L$-hop neighbors of $v$ via message passing, regardless of how large $G$ is. Hence, it suffices to explore and verify connected subgraphs in $G^{L}(V_{T})$ (see § II). In general, it interacts with three procedures:
@@
Auxiliary structures. $\mathsf{ASX}$-$\mathsf{OP}$ coordinates the interaction of the $\mathsf{Generator}$, $\mathsf{Verifier}$, and $\mathsf{Updater}$ via a state graph (simply denoted as $\zeta^{\prime}$). Each node (a “state”) $s\in\zeta^{\prime}$ records a candidate $G_{s}$ and its local information to be updated and used for evaluating ${\mathsf{SXQ}}^{k}$. There is a directed edge (a “transition”) $t$ = $(s,s^{\prime})$ in $\zeta^{\prime}$ if $G_{s^{\prime}}$ is obtained by applying a graph editing operator (e.g., edge insertion, edge deletion) to $G_{s}$. A path $\rho$ in the state graph $\zeta^{\prime}$ consists of a sequence of transitions that results in a candidate. In addition, each state $s$ is associated with (1) a score ${\sf DS}(G_{s})$, (2) a coordinate $\Phi(s)$, where each entry records an explainability measure $\phi(G_{s})$ ($\phi\in\Phi$); and (3) a variable-length bitvector $B(s)$, where an entry $B(s)[i]$ is $1$ if $G_{i}\preceq_{\epsilon}G_{s}$, and 0 otherwise. The vector $B(s)$ keeps the $\epsilon$-dominance relation between $G_{s}$ and current candidates in $\zeta^{\prime}$. Its ${\sf DS}(G_{s})$ score over $\zeta^{\prime}$, can be readily counted as the number of “1” entries in $B(s)$.
@@
3:set ${\mathcal{G}}_{\epsilon}$$\emptyset$;
4:identify edges for each hop: $\mathcal{E}$$\{E_{L},E_{L-1},\ldots,E_{1}\}$;
5:for $l$ = $L$ to 11 do
6:  initializes state $s_{0}$$G^{l}(v_{t})$.
@@
[…]
@@
8:  if $|{\mathcal{G}}_{\epsilon}|$<<$k$ then
@@
Update Dominance Relations. Procedure updateSX maintains a set of skyline explanations ${\mathcal{G}}_{s}$ (not shown) in terms of $(1+\epsilon)$-dominance. The set ${\mathcal{G}}_{s}$ is efficiently derived from the bitvectors $B(s)$ of the states in $\zeta^{\prime}$. The latter compactly encodes a lattice structure of $(1+\epsilon)$ dominance as a directed acyclic graph; and ${\mathcal{G}}_{s}$ refers to the states with no “parent” in the lattice, i.e., having an explanatory subgraph that is not $(1+\epsilon)$-dominated by any others so far. Details are in [47].
@@
Figure 2: Illustration of Onion Peeling (LL=22, kk=22). $\zeta^{\prime}$ = $\{s_{1},s_{2},s_{3},s_{5}\}$. $(\zeta^{\prime},\epsilon)$-explanation ${\mathcal{G}}_{\epsilon}$$\{s_{2},s_{5}\}$ with $\sf DS$} = $4$.
@@
Proof sketch: Consider procedure updateSX upon the arrival, at any time, of a new verified candidate $G_{s}$. (1) The above results clearly hold when $|\zeta^{\prime}|\leq k$ or $|{\mathcal{G}}_{\epsilon}|\leq k$, as ${\mathcal{G}}_{\epsilon}$ is the only $(\zeta^{\prime},\epsilon)$-explanation so far. (2) When $|{\mathcal{G}}_{\epsilon}|=k$, we reduce the approximate evaluation of ${\mathsf{SXQ}}^{k}$ to an instance of the online MAX $k$-SET problem [49]. The problem maintains a size-$k$ set cover with maximized weights. We show that updateSX adopts a greedy replacement policy by replacing a candidate in ${\mathcal{G}}_{\epsilon}$ with the new candidate $G_{s}$ only when this leads to a $(1+\frac{1}{k})$ factor improvement for $\sf DS$}. This ensures a $\frac{1}{4}$-approximation ratio [49]. Following the replacement policy consistently,  updateSX ensures a $\frac{1}{4}$ approximation ratio for $(\zeta^{\prime},\epsilon)$-explanations in $\zeta^{\prime}$. $\Box$
@@
[…]